Circular No. 8387

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004du IN UGC 11683 Further to IAUC 8386, D. Singer and W. Li report the LOSS/KAIT

discovery of an apparent supernova located at α = 21^{h}08^{m}26^{s}.51, δ = +18^{o}11′23′′.7 (equinox 2000.0), which is 3′′.4 west and 4′′.4 south of the center of UGC 11683. Unfiltered KAIT magnitudes: Aug. 9.29 UT, [20.0; 12.31, 18.4; 13.28, 17.1.

SUPERNOVA 2004dt IN NGC 799 M. Salvo, B. Schmidt, and P. Wood, Australian National University

(ANU), report that a spectrum (range 400–1000 nm; resolution 1.6 nm), taken with the ANU 2.3-m telescope (+ Double-Beam Spectrograph) at Siding Spring on Aug. 12.72 UT, shows SN 2004dt (cf. IAUC 8386) to be a young type-Ia supernova a few days before maximum light. The photospheric expansion velocity, deduced from the absorption minimum of Si II 635.5-nm, is ∼ 16500 km/s (via adoption of the NED recession velocity for the host galaxy NGC 799).

F. Patat and G. Pignata, European Southern Observatory; and S. Benetti, Istituto Nazionale di Astrofisica, on behalf of the ERTN, report that a fully reduced spectrum (range 320–850 nm; resolution 1.6 nm) of SN 2004dt (cf. IAUC 8386), obtained on Aug. 13.17 UT by J. Aceituno at the Calar Alto Observatory 2.2-m telescope (+ CAFOS), shows it to be a type-Ia supernova with some unusual features. In particular, there is a strong P-Cyg line whose absorption is measured at 392.4 nm (rest frame), tentatively identified as C II 426.7-nm, with an implied expansion velocity of ∼ 24100 km/s. If this identification is correct, the other intense C II 657.6–658.3-nm transitions are blended with the Si II 635.5-nm, which indeed appears very broad, intense, and with an asymmetric absorption trough. Particularly pronounced, with a broad symmetric P-Cyg profile, is also the O I 777.6-nm transition, from which an expansion velocity of ∼ 16800 km/s is deduced. In a normal type-Ia supernova, these rather high expansion velocities (the Ca II H and K line gives ∼ 22200 km/s) would indicate a very early phase. Other intense lines visible in the spectrum are from Fe III and Si III.

SUPERNOVA 2004cs IN UGC 11001 Further to IAUC 8386, the combination of B–g and R–V colors is

strongly suggestive of a type-Ia identification for this event (see Gal-Yam et al. 2004, PASP 116, 597).

2004 August 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green